I don't kow where to begin because area and perimeter are two seperate equations.. The area of a square is numerically equal to five times its perimeter. Find the length of a side of the square.
Answer by ankor@dixie-net.com(22740)

The area of a square is numerically equal to five times its perimeter.
Find the length of a side of the square.
:
let x = side of the square
:
Area = 5 times perimeter
x^2 = 5(4x)
x^2 = 20x
:
Subtract 20x from both sides
x^2 - 20x = 0
:
Factor out x
x(x-20) = 0
:
two solutions
x = 0
and the solution we want
x = 20 is side of the square
:
:
:
see if thats true
A = 20^2 = 400
P = 4(20) = 80, times 5 = 400 also